OpenStudy (anonymous):

Solve the equation using partial fractions, find an explicit solution if possible. Equation: dy/dx = (y^2-2y-3)dx

OpenStudy (anonymous):

\[y-3/y+1 = C \exp (4x)\]

OpenStudy (anonymous):

...so how did you get that?

OpenStudy (anonymous):

dy = (y-3)(y+1)dx dy/(y-3)(y+1) = dx 1/4(1/(y-3) -1/(y+1))dy = dx (partial fractions) integration will give you 1/4(ln(y-3) - ln(y+1)) = x + integration constant ln(y-3/y+1) = 4x + 4(integration constant) y-3/y+1 = e^(4x + K)= e^k (e^4x) y-3/y+1 = C e^4x

OpenStudy (anonymous):

okay awesome thanks :)